# LiDAR-BIND-T: Improved and Temporally Consistent Sensor Modality Translation and Fusion for Robotic Applications

### v2

@@ VI Experiments and Results
To evaluate the proposed improvements, we focus on both the reconstruction capabilities and SLAM performance, emphasising temporal consistency and improved temporal encoding by exploring and employing several metrics. For consistency and comparability, we use the same datasets as in our prior work. The experimental dataset can be summarised as follows: data was collected using a small robotic platform equipped with a 2D LiDAR (Hokuyo UST-20L) [27], mmwave radar (TI IWR1443) [28], and sonar (eRTIS) [6]. Sensor measurements were pre-processed to select the largest common field-of-view (FOV) for all sensors, which is an azimuth field of 100 degrees and a maximum range of approximately 5 meters. Recordings took place across multiple indoor environments, including hallways, offices and residential rooms. Our framework accommodates multiple datasets, with the mandatory condition that each contains the binding modality (LiDAR). Accordingly, the complete dataset consists of: (1) recordings in which all three sensors were used simultaneously; (2) several smaller sets covering subsets of these sensors; and (3) a simulated 2D LiDAR-only dataset employed for the initial stage of pre-training. This results in an average of 50,000 samples per sensor, and approximately 200,000 samples for LiDAR-only pre-training. Model training follows our earlier work, splitting each dataset into training and validation sets. For contrastive learning, larger batch sizes favour learning performance; thus, we utilise a batch size of 64 (constrained by available hardware).
@@
The FVMD score is calculated by first extracting motion trajectories of key points, typically using a pre-trained point tracking model such as PIPs++ [22]. Next, the Fréchet distance between the distributions of motion trajectories in the generated and reference video sets is calculated to obtain the final FVMD score. Lower FVMD scores indicate better motion consistency in the generated data. However, since the PIPs++ model does not support our specific data modality, we instead extract motion trajectories using a simple Lucas-Kanade optical flow algorithm, applied to a pre-defined grid in our heatmap images. Given these modifications, as well as the differences between our input data and typical usage scenarios, we cannot assume the validity of the FVMD score for our evaluation. Therefore, we further assess the metric’s effectiveness to verify its suitability for evaluating the temporal consistency of our models.
@@
  • SLAM performance: We evaluate the SLAM performance of the models by using the predicted LiDAR point clouds in a SLAM pipeline, specifically Google Carthographer [26]. The SLAM performance is evaluated using the Absolute Positional Error (APE), which measures the difference between the estimated trajectory and the ground truth trajectory. The ground truth trajectory is obtained from the wheel encoders of the robot. Next to that, we evaluate the occupancy maps match by calculating the IoU (Intersection over Union) between an L2L map and a map generated using the predicted point clouds from another model or fusion version described above. The maps are first aligned by finding the best transformation between the two maps using the cross-correlation between the two maps. We chose the L2L map as a reference here, as this map contains the amount of detail we can expect the SLAM algorithm to produce based on the model predictions. We show the SLAM performance metrics for two SLAM configurations: one where scan matching and odometry are used, and another where scan matching is the only source of transformation between the scans to obtain pose information.
@@
Next, the temporal transformer fusion model also demonstrates significant improvements in SLAM results, including lower APE and higher IoU, relative to the original LiDAR-BIND fusion model. The occupancy maps produced by the improved models are more consistent and exhibit greater alignment with the ground truth trajectory, as illustrated in Figure 10.
We also observe higher cosine similarity across the improved model versions compared to the original LiDAR-BIND models, indicating that the embeddings of the predicted frames are more similar to those produced by the L2L embeddings. This increased similarity should result in improved decoder performance, as embeddings are used as input to this module. The better aligned the predicted embeddings are with the L2L embeddings, the closer the predictions should match their LiDAR counterparts. The L2L model continues to offer the best overall performance across all metrics.
@@ VIII Discussion and Conclusion
Future work includes exploring the updated framework’s capabilities for single modality anomaly detection by tracking embedding orientation across time. Next to that, exploring attention-based and recurrent temporal modules for longer-range dependencies, and refining modality-specific encoders to further improve temporal stability and SLAM across diverse environments.

### v3

@@ VI Experiments and Results
To evaluate the proposed improvements, we focus on both the reconstruction capabilities and SLAM performance, emphasising temporal consistency and improved temporal encoding by exploring and employing several metrics. For consistency and comparability, we use the same datasets as in the prior LiDAR-BIND work. The experimental dataset can be summarised as follows: data was collected using a small robotic platform equipped with a 2D LiDAR (Hokuyo UST-20L) [28], mmwave radar (TI IWR1443) [29], and sonar (eRTIS) [7], Figure 9-b depicts the used robot setup. Sensor measurements were pre-processed to select the largest common field-of-view (FOV) for all sensors, which is an azimuth field of 100 degrees and a maximum range of approximately 5 meters. Recordings took place across multiple indoor environments, including hallways, offices and residential rooms. Our framework accommodates multiple datasets, with the mandatory condition that each contains the binding modality (LiDAR). Accordingly, the complete dataset consists of: (1) recordings in which all three sensors were used simultaneously; (2) several smaller sets covering subsets of these sensors; and (3) a simulated 2D LiDAR-only dataset employed for the initial stage of pre-training. This results in an average of 50,000 samples per sensor, and approximately 200,000 samples for LiDAR-only pre-training. Model training follows the earlier work, splitting each dataset into training and validation sets. For contrastive learning, larger batch sizes favour learning performance; thus, we utilise a batch size of 64 (constrained by available hardware).
Figure 9: (a) Example of processed range-azimuth heatmaps from the LiDAR-BIND dataset. These heatmaps represent processed sensor data, where for each azimuth angle, the range information is encoded in a 2D grid format. (b) The robotic platform used for data collection, equipped with a 2D LiDAR, mmWave radar, and sonar sensors.
@@
The FVMD score is calculated by first extracting motion trajectories of key points, typically using a pre-trained point tracking model such as PIPs++ [23]. Next, the Fréchet distance between the distributions of motion trajectories in the generated and reference video sets is calculated to obtain the final FVMD score. Lower FVMD scores indicate better motion consistency in the generated data. However, since the PIPs++ model does not support our specific data modality, we instead extract motion trajectories using a simple Lucas-Kanade optical flow algorithm, applied to a pre-defined grid in our heatmap images. Figure 11 presents example flow fields created using this method. We show the flow fields for the ground truth LiDAR data, as well as for the R2L and the temporal fusion model. These plots illustrate the differences in motion consistency between the different models, with the temporal fusion model showing a flow field more similar to the ground truth LiDAR data. Given these modifications, as well as the differences between our input data and typical usage scenarios, we cannot assume the validity of the FVMD score for our evaluation. Therefore, we further assess the metric’s effectiveness to verify its suitability for evaluating the temporal consistency of our models.
@@
  • SLAM performance: We evaluate the SLAM performance of the models by using the predicted LiDAR point clouds in a SLAM pipeline, specifically Google Carthographer [27]. The SLAM performance is evaluated using the Absolute Positional Error (APE), which measures the difference between the estimated trajectory and the ground truth trajectory. The ground truth trajectory is obtained from the wheel encoders of the robot. Furthermore, we evaluate the occupancy maps match by calculating the IoU (Intersection over Union) between an L2L map and a map generated using the predicted point clouds from another model or fusion version described above. The maps are first aligned by finding the best transformation between the two maps using the cross-correlation between the two maps. We chose the L2L map as a reference here, as this map contains the amount of detail we can expect the SLAM algorithm to produce based on the model predictions. We show the SLAM performance metrics for two SLAM configurations: one where scan matching and odometry are used, and another where scan matching is the only source of transformation between the scans to obtain pose information.
@@
Next, the temporal transformer fusion model also demonstrates significant improvements in SLAM results, including lower APE and higher IoU, relative to the original LiDAR-BIND fusion model. The occupancy maps produced by the improved models are more consistent and exhibit greater alignment with the ground truth trajectory, as illustrated in Figure 10. We also observe higher cosine similarity across the improved model versions compared to the original LiDAR-BIND models, indicating that the embeddings of the predicted frames are more similar to those produced by the L2L embeddings. This increased similarity should result in improved decoder performance, as embeddings are used as input to this module. The better aligned the predicted embeddings are with the L2L embeddings, the closer the predictions should match their LiDAR counterparts. The L2L model continues to offer the best overall performance across all metrics.
@@ VII Ablation study
Figure 11: Overview of the flow fields obtained using the Lucas-Kanade optical flow algorithm on a grid of points for different models. The leftmost column shows the ground truth LiDAR flow field, while the other columns show the flow fields obtained from different models. The flow fields are visualised by plotting the trajectory of the tracked points. The results illustrate that the proposed improvements lead to flow fields that more closely resemble the ground truth, indicating improved temporal consistency in the predictions.
@@ VIII Discussion and Conclusion
Future work includes exploring the updated framework’s capabilities for single modality anomaly detection by tracking embedding orientation across time. Next to that, exploring attention-based and recurrent temporal modules for longer-range dependencies, and refining modality-specific encoders to further improve temporal stability and SLAM across diverse environments. Finally, exploring dynamic windowing could be a promising strategy to deal with different sensor update rates and motion dynamics.